A circular coffee table has a diameter of 5 ft. What will it cost to have the top refinished if the company chares $3 per square foot for the refinishing

Found 2 solutions by edjones, unlockmath:
Answer by edjones(7311)

A=pi*r^2
pi*2.5^2=19.63...sq ft
19.63*3=$58.90
.
Ed

Answer by unlockmath(1122)

Hello,
Since we know the area of a circle is pi radius squared we can find the number of square feet by plugging in the numbers.
Area = 3.14 (2.5)^2 Do the math. (Note: It gives the diameter as 5 ft so the radius is 2.5 ft)
Area=19.625 square feet Now multiply $3 and we get:
Rounded off it comes to $58.88
RJ Toftness
www.math-unlock.com